OpenStudy (anonymous):

2 and 5/6 * 8 and 4/7= ?? Answer needs to be a mixed number in simplest terms

OpenStudy (phi):

First change the mixed numbers into improper fractions can you do that ?

OpenStudy (anonymous):

no

OpenStudy (phi):

2 5/6 (two and five-sixths) to change into an improper fraction multiply the 2 times the "bottom number" then add the top number can you try to do that ?

OpenStudy (anonymous):

17

OpenStudy (anonymous):

??

OpenStudy (phi):

and write the 17 over the 6 in other words 2 5/6 is the same as 17/6

OpenStudy (phi):

now change 8 4/7 to an improper fraction

OpenStudy (anonymous):

60

OpenStudy (phi):

8 times the bottom number then add the top number that answer becomes the top number of the improper fraction the bottom of the improper fraction is 7

OpenStudy (anonymous):

what

OpenStudy (anonymous):

so 60/7 thats all you had to say

OpenStudy (phi):

you got the 60 (for the top). remember the bottom stays 7

OpenStudy (anonymous):

is the answer 24 and 1/5

OpenStudy (phi):

now multiply two fractions \[ \frac{17}{6} \cdot \frac{60}{7} \] when multiplying it's always better to simplify first if yo can notice you have 60 up top and 6 below. 6 divides into 60 evenly you get \[ \frac{17}{1}\cdot \frac{10}{7} \]

OpenStudy (phi):

now multiply top times top and bottom times bottom \[ \frac{170}{7} \] now change back to a mixed number.

OpenStudy (phi):

divide 7 into 170. how many times does 7 go into 170? and what is the remainder?

OpenStudy (phi):